Last fall, a gardener planted 62 iris bulbs. She found that only 48 of the bulbs bloomed in the spring.

Find the empirical probability that an iris bulb of this type will bloom. Give answer as a fraction in lowest terms.
How many of the bulbs should she plant next fall if she would like at least 53 to bloom?

Answer by cleomenius(959)

The empirical probability that an iris bulb will bloom should be 48/62 or in lowest terms 24/31.
The number of bulbs to plant for 53 to bloom should be (x) * (24/31) = 53,
or 53 / (24/31) = 53 * (31/24) = 68.46.
Therefore, we must plant 69 bulbs to satisfy the probability that 53 will bloom.
Cleomneius.